# GenEDA: Unleashing Generative Reasoning on Netlist via Multimodal Encoder-Decoder Aligned Foundation Model

### v1

@@ Abstract
The success of foundation AI has motivated the research of circuit foundation models, which are customized to assist the integrated circuit (IC) design process. However, existing pre-trained circuit models are typically limited to standalone encoders for predictive tasks or decoders for generative tasks. These two model types are developed independently, operate on different circuit modalities, and reside in separate latent spaces, which restricts their ability to complement each other for more advanced applications. In this work, we present GenEDA, the first framework that aligns circuit encoders with decoders within a shared latent space. GenEDA bridges the gap between graph-based circuit representations and text-based large language models (LLMs), enabling communication between their respective latent spaces. To achieve the alignment, we propose two paradigms that support both open-source trainable LLMs and commercial frozen LLMs. Built on this aligned architecture, GenEDA enables three unprecedented generative reasoning tasks over netlists, where the model reversely generates the high-level functionality from low-level netlists in different granularities. These tasks extend traditional gate-type prediction to direct generation of full-circuit functionality. Experiments demonstrate that GenEDA significantly boosts advanced LLMs’ (e.g., GPT-4o and DeepSeek-V3) performance in all tasks.
@@ I Introduction
Circuit foundation model: two main types. As Fig. 1 shows, existing circuit foundation models can be categorized into two main types: (a) circuit encoder for predictive tasks, and (b) circuit decoder for generative tasks. TABLE I provides a detailed comparison of these two types of works. Circuit encoder refers to pre-trained models that encode circuits into general embeddings (i.e., circuit representation learning). Taking these embeddings as input, various lightweight downstream models can be fine-tuned for predictive design tasks, such as circuit functional reasoning [3, 4, 5, 6, 7] and circuit quality prediction [8, 9, 10, 11]. Circuit decoder refers to pre-trained models with circuit-related generative capability. Built mostly on large language models (LLMs), these decoders generate text outputs, such as circuits in RTL code [12, 13, 14], verification assertions [15, 16], EDA tool scripts [17], etc.
Limitation: lack of alignment. Developed independently, circuit encoders and decoders operate on different modalities and handle circuit representations in clearly distinct latent spaces. Specifically: (a) Circuit encoders typically work in the circuit graph latent space. They excel at capturing circuit structural and functional properties for predictive tasks, but lack generative capabilities. (b) Circuit decoders, typically LLMs, operate in the text latent space. They are effective at generating circuit-related text (e.g., RTL code, assertions), but they rely solely on textual input and cannot fully utilize the underlying structural information of circuits. As a result, these two important types are not aligned due to an inherent gap in their latent space, preventing more advanced capabilities.
Figure 1: (a) & (b) Two main types of existing circuit foundation models, including encoders for prediction and decoders for generation. (c) GenEDA proposes a general framework with alignment between circuit encoders and decoders.
GenEDA: an encoder-decoder alignment framework. In this work, we present GenEDA, the first framework that aligns pre-trained circuit encoders with LLM decoders. GenEDA bridges these two major types by communicating circuit graph latent space and text latent space, enabling effective information exchange. This alignment allows structural and functional insights captured by circuit encoders to directly enhance the generative capabilities of LLM decoders.
@@
Arch. Method Format Modality Pred. Gen. Description Direction
@@
MGVGA [9] AIG func./quality pred. R/F
Encoder NetTAG [19] Netlist Graph & Text func./quality pred. R/F
RTLCoder [20], etc. Spec RTL gen.
HDLDebugger [21], etc. RTL RTL debug
AssertLLM [15], etc Spec Text (Image) assertion gen. Forward
Decoder DeepRTL [22] RTL/Spec Text RTL understand./gen. R/F
Enc-Dec GenEDA Netlist Graph&Text Function gen. Reverse
@@
TABLE I: Comparison of GenEDA with representative categories of circuit foundation models. Existing circuit encoders mainly leverage graph structure for prediction tasks, while circuit decoders focus on semantic text for generation tasks. GenEDA bridges the gap between these two widely explored branches by aligning encoders and decoders within a shared latent space. After alignment, GenEDA supports more challenging functionality generation tasks.
Challenging application: generative reasoning on netlist. GenEDA’s alignment framework can support unprecedentedly challenging generative applications, such as generative reasoning on netlists [^1] In netlist functionality reasoning tasks, the high-level specification and RTL code as ground-truth are unknown to the model. Models are only provided with the low-level netlist as inputs.. Unlike RTL code, netlists are composed of a huge number of low-level, bit-blasted gates and their complex connections, offering little human-readable semantics for LLMs to understand. While prior netlist encoders [7, 23, 24, 10, 19] can extract structural and functional features into embeddings, they are limited to predictive tasks, only classifying the functionality of individual gates. GenEDA bridges this gap by aligning the encoder’s structural understanding of netlists with the decoder’s generative strengths. This encoder-decoder alignment enables generating high-level functionality directly from low-level netlist inputs, which is an unprecedentedly challenging task due to the irreversible nature of logic synthesis.
Specifically, GenEDA reasons functionalities of given netlists in a wide spectrum of granularities, with outputs including: (1) general function description, (2) circuit implementation details, and (3) fine-grained exact RTL code. These GenEDA-supported new generative netlist functional reasoning tasks are highly valuable in multiple aspects: (1) Practical applications: Reasoning high-level functionality from bit-blasted netlists can support critical applications such as functional verification, datapath optimization, and malicious logic detection. (2) Unprecedented reasoning quality: These tasks shift from traditional gate-level structural analysis to generating the overall functionality of entire circuits, including specification and RTL code, offering a significant leap in reasoning quality. (3) Benchmarking model capability: Our proposed tasks introduce new benchmarks for evaluating the generative capabilities and netlist understanding of foundation models. Since these tasks generate human-readable circuit information, they help enhance the interpretability of circuit models.
@@
  • Aligned circuit encoder-decoder framework. We propose the first framework that aligns pre-trained circuit encoders with LLM decoders for generative tasks. It supports both trainable and frozen decoders through two alignment paradigms.
  • Generative netlist foundation model. Building on this framework, we develop the first generative foundation model for netlists. By integrating structural and functional insights from netlist encoders, the model enhances LLM-based generation and enables reasoning over low-level, bit-blasted netlists.
  • New netlist generative reasoning tasks and benchmarks. We introduce three novel generative functional reasoning tasks on netlists, advancing beyond prior gate-type predictions. We also release corresponding benchmarks to encourage follow-up research on these tasks.
@@ II Related Work
Figure 2: Overview of GenEDA framework integrated into the standard IC design flow. GenEDA aligns the pre-trained netlist encoder with LLM decoders through two alignment paradigms, enabling challenging netlist generative functional reasoning tasks.
Figure 3: Proposed netlist generative reasoning benchmark. For Tasks 1 and 2, the netlist and question prompt are processed by models for specification reasoning, and evaluated using NLP scores and GPT scores. For Task 3, the models reconstruct RTL code from the netlist, which is evaluated via simulation and GPT scores.
@@ II-A Method Related: Circuit Foundation Model
Circuit encoders for prediction. Encoder-based models typically learn structured circuit representations to support predictive tasks such as reverse functional reasoning and early-stage design quality estimation. Most methods [6, 5, 4, 25, 18, 7, 24, 10, 26] focus on AIG netlists and use graph learning to capture the circuit structure. Recent work [8, 9, 19] fuses multiple modalities (graph, text, image), but their fusion methods are limited to the encoder side, lacking alignment with LLMs for better generation. Circuit decoders for generation. LLM-based decoders support forward-generation tasks like RTL or assertion generation [27, 20, 28, 15, 29], debugging [21], optimization [13, 30], and knowledge querying [31, 32]. DeepRTL [22] explores bidirectional generation between RTL and specification. However, RTL is easier for LLMs due to its rich semantics, making text-only models sufficient. Additionally, existing multimodal LLMs used in these tasks focus on visual inputs, without considering circuit graph structures. Circuit encoder-decoder alignment by GenEDA. GenEDA bridges this gap with the first encoder-decoder alignment framework, aligning representations in a shared latent space to support advanced generation tasks.
@@ II-B Application Related: Functional Reasoning on Netlists
Functional reasoning on gate-level netlists aims to reconstruct the high-level functionality originally described in specifications or RTL. It plays a critical role in functional verification, logic optimization, datapath synthesis, and hardware security. Existing approaches primarily fall into two categories: formal methods for analyzing functionality, and machine learning methods that focus on gate-level function prediction. We detail these two categories below.
Formal analysis. Traditional methods rely on structural and functional analysis using formal techniques [33, 34, 35]. These approaches typically extract subcircuits from the netlist and match them against components in a golden library via exhaustive formal verification. However, they are time-consuming, dependent on library completeness, and incapable of recognizing functional variants. GNN-based prediction. Recently, GNN-based methods [36, 37, 7, 23, 10] have been used for predictive netlist functional reasoning, focusing on gate-level function classification. While effective in identifying roles of known components, they require annotated labels and cannot generalize to unseen functionality or reason about the full circuit behavior.
Functionality generation by GenEDA. GenEDA moves beyond gate-level prediction to full-circuit generative reasoning. By aligning netlist encoders with LLM decoders, it enables direct generation of high-level specifications and RTL code from low-level gate-netlists, supporting semantics-aware understanding that goes far beyond previous prediction-based methods.
@@ III Overview
Fig. 2 presents the overview of our GenEDA framework, integrated into the standard digital IC design flow. GenEDA aligns the state-of-the-art post-synthesis netlist encoder NetTAG [19] $\mathbf{f}_{\text{enc}}$ with cutting-edge LLM-based decoders $\mathbf{f}_{\text{dec}}$. It first converts the circuit netlist $\mathcal{N}$ into embeddings via encoder $\mathbf{f}_{\text{enc}}$, capturing both netlist structural and functional information. The encoder output is then provided to aligned decoders $\mathbf{f}_{\text{dec}}$ to support advanced generative reasoning tasks on netlists.
@@
The rest of the paper is organized as follows: In Section IV, we first provide a detailed explanation of our three novel netlist generative reasoning tasks and our contributed benchmarks. In Section V, we elaborate the two proposed encoder-decoder alignment paradigms, as well as how to tackle the three tasks. Finally, in Section VI, we demonstrate the effectiveness of GenEDA in all three generative functionality reasoning tasks in experiments.
IV Netlist Generative Reasoning Tasks and Bechmarks
Fig. 3 illustrates our three novel generative reasoning tasks on netlists, along with the benchmarks developed to support them. These tasks aim to reversely generate high-level circuit functionality, including natural language specifications and exact RTL code, from low-level bit-blasted netlists. Please note that models are only provided with netlists, their corresponding specifications and RTL code serve as ground-truth and and unknown to the model. Our proposed benchmarks evaluate the generative model’s ability to truly understand the functionality of netlists, setting a new direction for generative EDA tasks. We detail the three tasks below.
Figure 4: Two encoder-decoder alignment paradigms in GenEDA: 1. Embedding-based alignment integrating embeddings from the netlist encoder with trainable LLM decoders, and 2. Prediction-based alignment using functional predictions from the netlist encoder as input for frozen LLM decoders.
@@ IV Netlist Generative Reasoning Tasks and Bechmarks
  • Question prompt. For Task 1, the prompt inquires the model to describe the interface, purpose, functionality, and constraints of the netlist. For Task 2, the prompt asks the model to explain the combinational logic, sequential behavior, and control flow.
  • Ground-truth specification text. Since real-world RTL specifications are rarely available, we generate reference answers for these two tasks using GPT-4o prompted with RTL code and manually verify their quality through expert review.
Evaluation metrics. As there are no standard metrics for evaluating the functional similarity of specification texts, we follow prior work DeepRTL [22], which addresses the RTL-to-specification task, and adopt a combination of natural language metrics. Specifically, we use BLEU, ROUGE-1/2/L, and cosine similarity computed from text embeddings. Additionally, GPT-4o is employed as an automated evaluator to assess semantic similarity between generated outputs and reference specifications. The prompt templates used are provided in Section VI.
@@
We propose the first benchmark for reverse RTL code reasoning. Given the task’s complexity, we begin by focusing on arithmetic blocks. Specifically, we extend the GNN-RE gate-level arithmetic function prediction dataset [36] into a generative benchmark, incorporating golden RTL code and testbenches. For each of the 9 arithmetic designs in our benchmark, we provide:
@@
  • Testbench. A verified testbench with predefined module name and IO ports, containing multiple input-output cases for functional validation.
Evaluation metrics. Similar to the existing forward RTL generation evaluation metrics [38, 39], we validate both syntax correctness and arithmetic functionality of the reversely generated RTL code using our provided golden testbenches. Additionally, GPT-4o is employed to assign a function similarity score, measuring how closely the generated RTL matches the ground truth. We provide detailed evaluation method implementation and prompts for obtaining GPT-score in Section VI.

### v2

# NetEncoder: General Netlist Representation Learning via Self-Supervised Text-Attributed Graph

@@ Abstract
The success of foundation AI has motivated the research of circuit foundation models, which are customized to assist the integrated circuit (IC) design process. However, existing pre-trained circuit foundation models are typically limited to standalone encoders for predictive tasks or decoders for generative tasks. These two model types are developed independently, operate on different circuit modalities, and reside in separate latent spaces. This restricts their ability to complement each other for more advanced capabilities. In this work, we present GenEDA, the first framework that cross-modally aligns circuit encoders with decoders within a shared latent space. GenEDA bridges the gap between graph-based circuit representation learning and text-based large language models (LLMs), enabling communication between their respective latent spaces. To achieve the alignment, we propose two paradigms to support both open-source trainable LLMs and commercial frozen LLMs. We leverage this aligned architecture to develop the first generative foundation model for netlists, unleashing LLMs’ generative reasoning capability on the low-level and bit-blasted netlists. GenEDA enables three unprecedented generative netlist functional reasoning tasks, where it reversely generates high-level functionalities such as specifications and RTL code from low-level netlists. These tasks move beyond traditional gate function classification to direct generation of full-circuit functionality. Experiments demonstrate that GenEDA significantly boosts advanced LLMs’ (e.g., GPT and DeepSeek series) performance in all tasks.
@@ I Introduction
Circuit foundation model: two main types. As Fig. 1 shows, existing circuit foundation models can be categorized into two main types: (a) circuit encoder for predictive tasks, and (b) circuit decoder for generative tasks. TABLE I provides a detailed comparison of these two types of works. Circuit encoder refers to pre-trained models that encode circuits into general embeddings (i.e., circuit representation learning). Taking these embeddings as input, lightweight downstream models are then fine-tuned for various predictive EDA tasks, such as circuit functional reasoning [3, 4, 5, 6, 7] and circuit quality prediction [8, 9, 10, 11]. Circuit decoder refers to pre-trained models with circuit-related generative capability. Built mostly on large language models (LLMs), these decoders generate text outputs, such as circuits in RTL code [12, 13, 14], verification assertions [15, 16], EDA tool scripts [17], etc.
Limitation: lack of alignment. Developed independently, circuit encoders and decoders operate on different modalities and handle circuit representations in clearly distinct latent spaces. Specifically: (a) Circuit encoders typically work in the circuit graph latent space. They excel at capturing circuit structural and functional properties into embeddings for predictive tasks, but lack generative capabilities. (b) Circuit decoders, typically LLMs, operate in the text latent space. They are effective at generating circuit-related text (e.g., RTL code, assertions), but they rely solely on textual input and cannot fully utilize the underlying structural information of circuits. As a result, these two important types are not aligned due to an inherent gap in their latent space, preventing more advanced capabilities.
Figure 1: (a) & (b) Two main types of existing circuit foundation models: encoders for prediction and decoders for generation. (c) GenEDA proposes a general framework that aligns circuit encoders with decoders, leveraging encoder-captured information to enhance decoder generation.
GenEDA: an encoder-decoder alignment framework. To address this, we present GenEDA [^1] The code of GenEDA and benchmarks of proposed tasks are available at: https://github.com/hkust-zhiyao/GenEDA, the first framework that aligns pre-trained circuit encoders with LLM decoders. GenEDA bridges these two major types by communicating circuit graph latent space and text latent space, enabling effective information exchange. This alignment allows structural and functional insights captured by circuit encoders to directly enhance the generative capabilities of LLM decoders.
@@
Type Method Format Modality Pred. Gen. Description Direction
@@
MGVGA [9] AIG func./quality pred. Reverse/Forward
Encoder NetTAG [19] Netlist
Graph
& Text
func./quality pred. Reverse/Forward
RTLCoder [20], etc. Spec RTL gen. Forward
HDLDebugger [21], etc. RTL RTL debug Forward
AssertLLM [15], etc Spec
Text
assertion gen. Forward
Decoder DeepRTL [22] RTL/Spec Text RTL understand./gen. Reverse/Forward
Enc-Dec GenEDA (ours) Netlist Graph&Text Function gen. Reverse
  • We list one of the representative works for each method in the table. For a more comprehensive comparison, please refer to Section II-A.
@@
TABLE I: Comparison of GenEDA with representative categories of circuit foundation models. Existing circuit encoders mainly leverage graph structure for prediction tasks, while circuit decoders focus on semantic text for generation tasks. GenEDA bridges the gap between these two widely explored branches by aligning encoders and decoders within a shared latent space. After alignment, GenEDA supports more challenging generative netlist functional reasoning tasks.
Challenging application: generative functional reasoning of netlist. GenEDA’s alignment framework can support unprecedentedly challenging generative applications, such as generative reasoning of the netlist functionality [^2] In the reverse netlist functional reasoning scenario, the high-level specification and RTL code as ground-truth are unknown to the model. Models are only provided with the low-level netlist as inputs.. Unlike RTL code, netlists are composed of a huge number of low-level, bit-blasted gates and their complex connections, offering little human-readable semantics for LLMs to understand. While prior netlist encoders [7, 23, 24, 10, 19] can extract structural and functional features into embeddings, they are limited to predictive reasoning, only classifying the functionality of individual gates. GenEDA bridges this gap by aligning the encoder’s structural and functional understanding of netlists with the decoder’s generative strengths. This encoder-decoder alignment enables generating high-level functionality directly from low-level netlist inputs, which is an unprecedentedly challenging task due to the irreversible nature of logic synthesis.
Specifically, GenEDA reasons functionalities of given netlists in a wide spectrum of granularities, with outputs including: (1) general function description, (2) circuit implementation details, and (3) fine-grained exact RTL code. These GenEDA-supported new generative netlist functional reasoning tasks are highly valuable in multiple aspects: (1) Practical applications: Reasoning high-level functionality from bit-blasted netlists can support critical applications [25, 26, 23], such as functional verification, datapath optimization, and malicious logic detection. (2) Unprecedented reasoning quality: These tasks move beyond traditional gate function classification by directly generating the overall functionality of entire circuits, including specifications and RTL code, offering a significant leap in reasoning quality. (3) Benchmarking model capability: Our proposed tasks introduce new benchmarks for evaluating the generative circuit reasoning capabilities and netlist understanding of foundation models. Since these tasks generate human-readable circuit information, they help enhance the interpretability of circuit models.
@@
  • Circuit encoder-decoder alignment framework. We propose the first framework that cross-modally aligns pre-trained circuit encoders with LLM decoders. It supports both trainable and frozen LLMs for advanced generative tasks through two alignment paradigms.
  • Generative netlist foundation model. Leveraging this framework, we develop the first generative foundation model for netlists. By integrating structural and functional insights captured by netlist encoders, the model unleashes LLMs’ generative reasoning capability on the low-level and bit-blasted netlists.
  • New generative netlist reasoning tasks and benchmarks. We propose three novel generative netlist functional reasoning tasks with corresponding benchmarks, advancing beyond prior gate function classification. We also release these benchmarks to encourage follow-up research on these tasks.
@@ II Related Work
Figure 2: Overview of GenEDA framework integrated into the standard IC design flow. GenEDA aligns the pre-trained netlist encoder with LLM decoders through two alignment paradigms, enabling challenging generative netlist functional reasoning tasks.
Figure 3: Proposed generative netlist functional reasoning benchmarks. For Tasks 1 and 2, the netlist and question prompt are processed by models for specification reasoning, and evaluated using NLP scores and GPT scores. For Task 3, the models reconstruct RTL code from the netlist, which is evaluated via simulation and GPT scores.
@@ II-A Method Related: Circuit Foundation Model
Circuit encoders for prediction. Encoder-based models typically learn structured circuit representations to support predictive tasks such as reverse functional reasoning and early-stage design quality estimation. Most methods [6, 5, 4, 27, 18, 7, 24, 10, 28] focus on AIG netlists and use graph learning to capture the circuit structure. Recent work [8, 9, 19] fuses multiple modalities (graph, text, image), but their multimodal fusion are limited to the encoder side, lacking alignment with LLMs for enabling generation capabilities. Circuit decoders for generation. LLM-based decoders support forward-generation tasks like RTL or assertion generation [29, 20, 30, 15, 31], debugging [21], optimization [13, 32], and knowledge querying [33, 34]. DeepRTL [22] further explores bidirectional generation between RTL and specification. However, since RTL is easier for LLMs to understand due to its rich semantics, the text-only T5 LLM is sufficient for these tasks. Additionally, existing multimodal approaches focus on circuit images [32, 15], a capabiltity already present in current multimodal LLMs. They overlook the crucial circuit graph structure, which remains largely unexplored in combination with LLMs.
Circuit encoder-decoder alignment by GenEDA. GenEDA bridges this gap by proposing the first cross-modal encoder-decoder alignment framework, which aligns circuit structural representations with LLM textual representations in a shared latent space, enabling support for advanced generative tasks.
@@ II-B Application Related: Functional Reasoning on Netlists
Functional reasoning on gate-level netlists aims to reconstruct the high-level functionality originally described in specifications or RTL. It plays a critical role in functional verification, logic optimization, datapath synthesis, and hardware security. Existing approaches primarily fall into two categories: formal methods for analyzing functionality [25, 35, 26, 36], and machine learning methods for gate-level function classification [37, 38, 7, 23, 10]. We detail these two categories below.
Formal analysis. Traditionally, netlist functional reasoning relies on structural and functional analysis using formal techniques [25, 35, 26, 36]. These approaches typically extract subcircuits from the netlist and match them against components in a golden library via exhaustive formal verification. However, they are time-consuming, dependent on library completeness, and incapable of recognizing functional variants. GNN-based individual gate function prediction. Recently, GNN-based methods [37, 38, 7, 23, 10] have been applied to predictive netlist functional reasoning. These methods focus on learning the netlist structure for gate-level function classification. While they effectively identify the roles of known components, they are limited in generalizing to unseen functionality and cannot reason about the full behavior of the entire circuit.
GenEDA-based entire circuit function generation. GenEDA advances from gate-level prediction to full-circuit generative reasoning. By aligning netlist encoders with LLM decoders, it enables the direct generation of high-level specifications and RTL code from low-level netlists.
@@ III Overview
Fig. 2 presents the overview of our GenEDA framework, integrated into the standard digital IC design flow. GenEDA aligns the state-of-the-art post-synthesis netlist encoder NetTAG [19] $\mathbf{f}_{\text{enc}}$ with cutting-edge LLM-based decoders $\mathbf{f}_{\text{dec}}$. It first converts the circuit netlist $\mathcal{N}$ into embeddings via the encoder $\mathbf{f}_{\text{enc}}$, capturing both netlist structural and functional information. The encoder output is then provided to aligned decoders $\mathbf{f}_{\text{dec}}$ to support advanced generative reasoning tasks on netlists.
@@
The rest of the paper is organized as follows: In Section IV, we first provide a detailed explanation of our three novel generative netlist functional reasoning tasks and our contributed benchmarks. In Section V, we illustrate the two proposed encoder-decoder alignment paradigms, as well as how to tackle the three tasks. In Section VI, we demonstrate the effectiveness of GenEDA in all three generative functional reasoning tasks in experiments. Finally, in Section VII, we discuss the potential of extending GenEDA to other circuit design stages and the broader applications of generative netlist functional reasoning.
IV Benchmarking Generative Functional Reasoning
Fig. 3 illustrates our three novel generative functional reasoning tasks on netlists, along with the benchmarks developed to support them. These tasks aim to reversely generate high-level circuit functionality, including natural language specifications and exact RTL code, from low-level bit-blasted netlists. Please note that models are only provided with netlists, their corresponding specifications and RTL code serve as ground-truths and and unknown to the model. Our proposed benchmarks evaluate the generative model’s ability to truly understand the functionality of netlists, setting a new direction for generative EDA tasks. We detail the three tasks below.
Figure 4: Two encoder-decoder alignment paradigms in GenEDA: 1. Embedding-based alignment integrating embeddings from the netlist encoder with trainable LLM decoders, and 2. Prediction-based alignment using functional predictions from the netlist encoder as textual input for frozen LLM decoders.
@@ IV Benchmarking Generative Functional Reasoning
  • Question prompt. For Task 1, the prompt inquires models to describe the interface, purpose, functionality, and constraints of the netlist. For Task 2, the prompt asks the model to explain the combinational logic, sequential behavior, and control flow.
  • Ground-truth specification text. Since real-world RTL specifications are rarely available, following [22, 8], we generate ground-truth reference answers for these two tasks using GPT-4o prompted with RTL code and manually verify their quality through expert review.
Evaluation metrics. As there are no standard metrics for evaluating the functional similarity of specification texts, we follow prior work DeepRTL [22] (i.e., generating specification from RTL code) to adopt a combination of natural language metrics. Specifically, we use BLEU, ROUGE-1/2/L, and cosine similarity computed from text embeddings. Additionally, GPT-4o is employed as an automated evaluator to assess semantic similarity between generated outputs and reference specifications. The prompt templates used are provided in Section VI.
@@
We propose the first benchmark for reverse RTL code reasoning. Given the task’s complexity, we begin by focusing on arithmetic modules. Specifically, we extend the GNN-RE gate-level arithmetic function prediction dataset [37] into a generative benchmark, incorporating golden RTL code and testbenches. For each of the 9 arithmetic designs in our benchmark, we provide:
@@
  • Testbench. A verified testbench with pre-defined module name and IO ports, containing multiple input-output cases for functional validation.
Evaluation metrics. Similar to the existing forward RTL generation evaluation metrics [39, 40], we validate both syntax correctness and function correctness of the reversely generated arithmetic RTL code using our provided golden testbenches. Additionally, GPT-4o is employed to assign a function similarity score, measuring how closely the generated RTL matches the ground truth. We provide detailed evaluation method implementation and prompts for obtaining GPT-score in Section VI.